The length of a rectangular swimming pool is twice its width. The pool is surrounded by a cement walk 4 feet wide. If the area of the walk is 748 feet^2, determine the dimensions of the pool.: The length of a rectangular swimming pool is twice its width. The pool is surrounded by a cement walk 4 feet wide. If the area of the walk is 748 feet^2, determine the dimensions of the pool.
Answer by ankor@dixie-net.com(4538)

the length of a rectangular swimming pool is twice its width. The pool is surrounded by a cement walk 4 feet wide. If the area of the walk is 748 ft^2, determine the dimensions of the pool.
:
Let x = the width of the pool
then
2x = the length
and
2x^2 = area of the pool
:
Overall area including the walk = (x+8)*(2x+8) = 2x^2 + 24x + 64
:
The equation:
Overall area - pool area = 748 sq/ft (walkway area)
(2x^2 + 24x + 64) - 2x^2 = 748
:
24x = 748 - 64
:
24x = 684
x = 684/24
x = 28.5 ft width of the pool
and
2(28.5) = 57 ft, length of the pool
:
:
Check solution:
Overall area = 65 * 36.5 = 2372.5
---Pool area = 57 * 28.5 = 1624.5
----------------------------------
leaving the walkway area = 748